# Explanation Space: A New Perspective into Time Series Interpretability

### v1

@@ Abstract
Human understandable explanation of deep learning models is necessary for many critical and sensitive applications. Unlike image or tabular data where the importance of each input feature (for the classifier’s decision) can be directly projected into the input, time series distinguishable features (e.g. dominant frequency) are often hard to manifest in time domain for a user to easily understand. Moreover, most explanation methods require a baseline value as an indication of the absence of any feature. However, the notion of lack of feature, which is often defined as black pixels for vision tasks or zero/mean values for tabular data, is not well-defined in time series. Despite the adoption of explainable AI methods (XAI) from tabular and vision domain into time series domain, these differences limit the application of these XAI methods in practice. In this paper, we propose a simple yet effective method that allows a model originally trained on time domain to be interpreted in other explanation spaces using existing methods. We suggest four explanation spaces that each can potentially alleviate these issues in certain types of time series. Our method can be readily adopted in existing platforms without any change to trained models or XAI methods. The code is posted in supplementary material.
Introduction
Explainable AI (XAI) is a recent branch of studies dedicated to provide human interpretable explanation of deep models. Initially introduced for image data, attribution-based methods aim to highlight the region of the image responsible for the model’s decision. Similar approach has been later adopted for other data types, including time series. However, unlike a typical image, time series are vastly diverse in nature. In many cases, the associated feature cannot be easily presented in time domain (Schröder, Zamanian, and Ahmidi 2023), such as dominant frequency, Hurst parameter, Lyapunov exponent, etc. In such cases, a classifier can achieve high accuracy while XAI methods are inevitably doomed to provide understandable explanation. For instance, the representation in time versus frequency domain of a model trained on FordA dataset is shown in Figure 1a and 1b which clearly shows that only a few frequency components are dominant, and thus more desirable, while the time domain explanation is more complex and hard to comprehend. Similarly, an audio signal the time/frequency space (Figure 1d) generates a more sparse explanation than the time domain (Figure 1c).
@@ Introduction
In this paper, our premise is that a more sparse explanation is more interpretable to an end user. Our key observation is that different spaces generate explanation with different sparsity on different type of time series that is aligned with our intuitive understanding of the nature of time series. The main challenge is that how to explain an existing DL model trained only on time domain in different domains.
@@
We show that the explanation spaces are not limited to the well-known time and frequency spaces. In fact, any one-to-one function can introduce a new explanation space. To show the usefulness of this generalization in practice, we introduce two new explanation spaces, namely, min-zero and difference spaces, each suitable for certain types of time-series, and addressing the baseline problem. We evaluate our method on nine XAI methods using three DL models and different types of time series, including sensor data, audio, and motion, and illustrate the usefulness of different spaces in explaining different types of time series with less complexity.
Related Work
@@
This paper focuses on post-hoc explanation methods. There are several mechanisms to generate an explanation: Using gradient, such as Integrated Gradient (Sundararajan, Taly, and Yan 2017), GradientSHAP (Erion et al. 2020), DeepLIFT (Shrikumar, Greenside, and Kundaje 2017), DeepSHAP (Lundberg and Lee 2017), Saliency (Simonyan, Vedaldi, and Zisserman 2013), InputXGradient (Shrikumar et al. 2016), Guided BackProp (Springenberg et al. 2014), and KernelSHAP (Lundberg and Lee 2017), or occluding features, such as Feature Occlusion (Zeiler and Fergus 2014), and feature ablation (Suresh et al. 2017). Most early works were originally developed for computer vision. Recently, some explanation methods are proposed to deal with challenges in time series data, such as TSR (Ismail et al. 2020), Demux (Doddaiah et al. 2022), TimeX (Queen et al. 2024), LEFTIST (Guillemé et al. 2019), MILLET (Early et al. 2023), or DFT-LPR (Vielhaben et al. 2024). All these methods, and majority of computer vision methods, are attribution-based highlighting the importance of each pixel/time-step to the prediction.
@@
In this paper, we focus on a new concept, i.e. explanation space, rather than a new explanation method. The closest work in literature is DFT-LPR (Vielhaben et al. 2024) where the authors proposed a virtual inspection layer at the beginning of a model trained on time domain. They adopted LPR (Bach et al. 2015) method such that it propagates the relevance back to the frequency or time/frequency domain. However, this method only works with LPR and also doesn’t easily allow new spaces to be introduced. In this paper, we solve this limitation by introducing a generalized wrapping mechanism that changes the input type of a model without changing its functionality. Hence, any existing XAI method can be used to generate explanation in the new input type.
Method
@@
Let $x=\{x_{1},x_{2},...,x_{N}\}\in R^{N}$ be a time series sample, where $x_{i}$ is the value at time step $i$ and $N$ is the number of time steps. To each time series sample there is a label associated, denoted by $y\in\{y_{1},...,y_{C}\}$ where $C$ is the number of classes. A classifier, $M:x\rightarrow y$, maps a times series sample to a probability distribution vector over $C$ classes. An explanation method, $E(M(.),x)\rightarrow\bar{x}$, takes a classifier and a time series sample and generates an output, $\bar{x}$. In the case of attribution methods, $\bar{x}$ is a heat-map with the same dimension as $x$, and in the case of counterfactual methods, $\bar{x}$ is a new time series where $M(x)\neq M(\bar{x})$.
@@
Projection of Explanation from Time Domain into a New Explanation Space: Assume a classifier $M$ trained on time domain. Consider an explanation space $F$ (e.g. FFT where it outputs frequency domain). Since $M(F^{-1}(F(x)))=M(x)$, we can define $z=F(x)$ as a new domain and the associated classifier is $M^{\prime}(z)=M(F^{-1}(z))\rightarrow y$. The new classifier $M^{\prime}$ is the same as $M$ except that it performs the reverse operation, $F^{-1}$, before passing the input to $M$. Hence, $M$ on $x$ domain is functionally the same as $M^{\prime}$ on $z$. However, now, we can use any existing explanation method, $E$, to generate an explanation on $z$ domain (or on explanation space $F$) using $E(M^{\prime}(.),z)$ [^2] It is desirable for $F/F^{-1}$ to be almost everywhere differentiable. Otherwise, gradient-based explanation methods may fail..
Explanation Spaces
@@
Time/Frequency Space is the short-time discrete Fourier transform to allow the explanation on both time and frequency domains. Figure 1c shows the attribution generated by DeepLIFT on AudioMNIST (Becker et al. 2023) in time domain and Figure 1d shows the attribution generated by the same method on time/frequency domain. The attribution on time/frequency domain is more focused on the center of the time series where actual signal is presented.
@@
Min Zero Space aims to alleviate the well-known issue of baseline exist on many attribution-based methods (Höllig, Thoma, and Grimm 2023). By default, many existing explanation methods implicitly or explicitly assume a specific value (zero for the most part or a value set by user or training set) to be associated with a lack of any feature, including InputXGradient, feature occlusion, DeepLIFT, etc. Because time series are often z-normalized per sample to have a zero mean mainly for training stability, the zero value does not correspond to the lack of feature even if the original unnormalized time series does. For example, power consumption time series (e.g. ElectricDevices and LargeKitchenAppliances datasets in UCR repository) or time series dataset containing number of events or magnitude of events (e.g. Earthquakes and ChinaTown datasets in UCR repository or network traffic classification (Rezaei and Liu 2020)) are examples of natural time series where zero corresponds to a lack of feature/event in the unnormalized version, but not in the normalized one. This causes many artifacts to appear in XAI methods as shown in Figure 2a. This is consistent with the observation in (Samek et al. 2019) that a global mean shift in data changes the explanation.
@@
where $x_{min}$ is the minimum value in the normalized time series. Note that the reverse operation ($F^{-1}$) is needed to be embedded into a new neural network, $M(F^{-1}(.))$. This can be easily done by adding a single fully connected layer before $M()$ where it adds the last element to the first N elements. The implementation of the weight matrix associated with such a fully connected layer is explained in Appendix. As shown in Figure 2b, generating the attribution on mean zero space using the same method as Figure 2a completely removes all artifacts assigned to the part with no signal.
@@
Difference Space is a result of taking a difference of each two consecutive time steps as a new time series. Differencing has been widely used in time series analysis to remove a trend or to make a time series stationary (Hyndman and Athanasopoulos 2018). In our framework, difference domain has two benefits. First, for non-stationary time series (often with trend), such as stock price or atmospheric time series, a shapelet-like feature in a small region can be completely imperceptible. For example, in a synthetic data generated to mimic a non-stationary time series with a small shapelet as a distinguishable feature, a shapelet feature may not be easily visible when it is small in comparison to the range of the entire time series [^3] We use synthetic data in this experiment since there is no dataset with such properties in UCR repository or any publicly available classification repository we are aware of., as shown in Figure 3a and 3c. However, the shapelets (i.e. semi-linear increase/decrease patterns visible in difference domain) can be clearly located in difference domain where the long range fluctuations is effectively removed, as shown in the corresponding samples in Figure 3b and 3d. Interestingly, the difference and time domain has a huge impact on the performance of XAI methods. The XAI method completely fails the time domain which can be explained by the second benefit.
Second, the difference domain implicitly changes the notion of a feature. Here, the lack of change in time domain corresponds to zeros in difference domain (implying a lack of feature), and changes correspond to non-zeros values. This can solve the baseline issue for certain typres of time series. For example, in time series associated with movements (e.g. GunPoint) or ECG signals, the lack of activity is better manifested by a sequence of near constant values, not necessarily a sequences of zeros. By taking the difference of the time series, a sequence of unchanging values will map to a sequence of zeros and, hence, aligns with the most explanation methods’ baselines. Formally, the difference space is defined as
@@
The reverse operation is trivial and the fully connected implementation is explained in Appendix. Figure 2c and 2d shows the attribution generated by DeepLIFT on time domain and difference domain, respectively. Note that in Figure 2d we show the attribution generated on difference space on top of the time domain just for better visualization. In reality, the difference domain time series looks different from the original time domain. As shown, the difference space’s explanation is sparser and focuses on the actual heartbeat of the ECG signal.
Evaluation
@@
Experimental Setup: We include several datasets covering different types of time series, as shown in Table 1, mainly from UCR Repository (Dau et al. 2019) except AudioMNIST provided in (Vielhaben et al. 2024). For each dataset, we trained two commonly used model architectures: a ResNet [^4] ResNet implementation for time series is adopted from TSEvo package (Höllig, Kulbach, and Thoma 2022). and an InceptionTime model (Ismail Fawaz et al. 2020). We emphasize that all models are trained only on time domain. The explanations are generated on different domain using the process described in Method Section. We evaluate nine well-known XAI methods: DeepLIFT, GradientSHAP, Guided BackProp (Springenberg et al. 2014), InputXGradient (I×$\times$×G), Integrated Gradient (IG), KernelSHAP, LIME (Ribeiro, Singh, and Guestrin 2016), Occlusion, and Saliency (Simonyan, Vedaldi, and Zisserman 2013). We use Captum implementation (Kokhlikyan et al. 2020) for XAI methods.
@@
Note that the goal of this paper is not to have comprehensive evaluation of different XAI methods, as it has previously shown that no single XAI method can outperform others in all metrics/datasets (Löffler et al. 2022). Rather, we focus on illustrating that the concept of spaces is useful and that the same XAI method may bring a better explanation (with respect to some metric such as sparsity) on different spaces depending on the type of time series. While our method can be easily adopted for existing counter-factual XAI methods as well, such evaluation is the subject of future studies.
@@
Table 1: Dataset types and models’ accuracy
Evaluation Criteria
Robustness: Robustness of an attribution method refers to its sensitivity with respect to small perturbations of the target sample (Höllig, Thoma, and Grimm 2023; Löffler et al. 2022). Unlike previous studies, we investigate the effect of different input spaces. As a result, a trained model wrapped to take frequency domain input may be more sensitive than the unwrapped version taking the time domain input. It is not trivial whether a model wrapped in other domains is more sensitive to perturbation than time domain or not. Hence, we also need to evaluate the sensitivity of the model itself in different input spaces. So, unlike previous work, we define two robustness: the first one is associated with the classifier and the second one is associated with the XAI method. Classifier robustness is defined as follow:

### v3

@@ Abstract
Human understandable explanation of deep learning models is essential for various critical and sensitive applications. Unlike image or tabular data where the importance of each input feature (for the classifier’s decision) can be directly projected into the input, time series distinguishable features (e.g. dominant frequency) are often hard to manifest in time domain for a user to easily understand. Additionally, most explanation methods require a baseline value as an indication of the absence of any feature. However, the notion of lack of feature, which is often defined as black pixels for vision tasks or zero/mean values for tabular data, is not well-defined in time series. Despite the adoption of explainable AI methods (XAI) from tabular and vision domain into time series domain, these differences limit the application of these XAI methods in practice. In this paper, we propose a simple yet effective method that allows a model originally trained on the time domain to be interpreted in other explanation spaces using existing methods. We suggest five explanation spaces, each of which can potentially alleviate these issues in certain types of time series. Our method can be easily integrated into existing platforms without any changes to trained models or XAI methods. The code will be released upon acceptance.
Index Terms:
XAI, uncertainty principle, time series.
I Introduction
Explainable AI (XAI) is a recent branch of studies dedicated to providing human interpretable explanation of deep models. Initially introduced for image data, attribution-based methods aim to highlight the region of the image responsible for the model’s decision. Similar approaches have been later adopted for other data types, including time series. However, unlike computer vision, time series is an umbrella term covering vastly different applications with widely different characteristics and natures. This includes stock market data, EEG signal, audio signals, accelerometer data, as well as non-temporal data, such as spectrogram (e.g. Beef and Coffee dataset in UCR repository) or skeleton(contour)-based descriptors (e.g. ArrowHead or Fish datasets). Although application-blind DL approaches have often worked largely for classifications due to the power of DL, we face significant challenges when it comes to explanation.
For instance, the explanation in time versus frequency domain of a model trained on FordA dataset is shown in Figure 1a and 1b which clearly shows that only a few frequency components are dominant, and thus more desirable, while the time domain explanation is more complex and hard to comprehend. Similarly, in audio signal the time/frequency space (Figure 1d) generates a more sparse explanation than the time domain (Figure 1c).
@@ I Introduction
In this paper, our premise is that a more sparse explanation is more interpretable to an end user, and more likely to reveal the nature of the time series. Our key observation is that different spaces generate explanation with different sparsity on different type of time series that is aligned with our intuitive understanding of the nature of time series. More importantly, different spaces can reveal different aspects of data generation process, e.g. shapelet-based features in time domain and wave-based features in frequency. The main challenge is how to explain an existing DL model trained only on time domain in different domains.
@@
We show that the explanation spaces are not limited to the well-known time and frequency spaces. In fact, any one-to-one function can introduce a new explanation space. To show the usefulness of this generalization in practice, we introduce three new explanation spaces that has not been used in XAI literature, namely, min-zero, difference, and decomposition spaces, each suitable for certain types of time-series. We evaluate our method on nine XAI methods using two DL models and different types of time series, including sensor data, audio, and motion, and illustrate the usefulness of different spaces in explaining different types of time series with less complexity.
II Related Work
@@
This paper focuses on post-hoc explanation methods. There are several mechanisms to generate an explanation: Using gradient, such as Integrated Gradient [1], GradientSHAP [2], DeepLIFT [3], DeepSHAP [4], Saliency [5], InputXGradient [6], Guided BackProp [7], and KernelSHAP [4], or occluding features, such as Feature Occlusion [8], and feature ablation [9]. Most early works were originally developed for computer vision. Recently, some explanation methods are proposed to deal with challenges in time series data, such as TSR [10], Demux [11], TimeX [12], LEFTIST [13], MILLET [14], or DFT-LPR [15]. All these methods, and majority of computer vision methods, are attribution-based highlighting the importance of each pixel/time-step to the prediction.
@@
In this paper, we focus on a new concept, i.e. explanation space, rather than a new explanation method. The closest work in literature is DFT-LPR [15] where the authors proposed a virtual inspection layer at the beginning of a model trained on time domain. They adopted LPR [16] method such that it propagates the relevance back to the frequency or time/frequency domain. However, this method only works with LPR and also doesn’t easily allow new spaces to be introduced. In this paper, we address this limitation by introducing a generalized wrapping mechanism that changes the input type of a model without changing its functionality. Hence, any existing XAI method can be used to generate explanation in the new input type. We show that the complexity introduced in the inspection layer of LPR is unnecessary and a simple, yet generalizable, approach can be adopted.
III Method
@@
Let $x=\{x_{1},x_{2},...,x_{N}\}\in R^{N}$ be a time series sample, where $x_{i}$ is the value at time step $i$ and $N$ is the number of time steps. To each time series sample there is a label associated, denoted by $y\in\{y_{1},...,y_{C}\}$ where $C$ is the number of classes. A classifier, $M:x\rightarrow y$, maps a times series sample to a probability distribution vector over $C$ classes. An explanation method, $E(M(\cdot),x)\rightarrow\bar{x}$, takes a classifier and a time series sample and generates an output, $\bar{x}$. In the case of attribution methods, $\bar{x}$ is a heat-map with the same dimension as $x$, and in the case of counterfactual methods, $\bar{x}$ is a new time series where $M(x)\neq M(\bar{x})$.
@@
Projection of Explanation from Time Domain into a New Explanation Space: Assume a classifier $M$ trained on time domain. Consider an explanation space $F$ (e.g. FFT where it outputs frequency domain). Since $M(F^{-1}(F(x)))=M(x)$, we can define $z=F(x)$ as a new domain and the associated classifier is $M^{\prime}(z)=M(F^{-1}(z))\rightarrow y$. The new classifier $M^{\prime}$ is the same as $M$ except that it performs the reverse operation, $F^{-1}$, before passing the input to $M$. Hence, $M$ on $x$ domain is functionally the same as $M^{\prime}$ on $z$. However, now, we can use any existing explanation method, $E$, to generate an explanation on $z$ domain (or on explanation space $F$) using $E(M^{\prime}(\cdot),z)$ [^2] It is desirable for $F/F^{-1}$ to be almost everywhere differentiable. Otherwise, gradient-based explanation methods may fail..
III-A Explanation Spaces
@@
Time/Frequency Space is the short-time discrete Fourier transform to allow the explanation on both time and frequency domains. Figure 1c shows the attribution generated by DeepLIFT on AudioMNIST [17] in time domain and Figure 1d shows the attribution generated by the same method on time/frequency domain. The attribution on time/frequency domain is more focused on the center of the time series where actual signal is presented.
@@
Min Zero Space aims to alleviate the well-known issue of baseline exist on many attribution-based methods [18]. By default, many existing explanation methods implicitly or explicitly assume a specific value (zero for the most part or a value set by user or training set) to be associated with a lack of any feature, including InputXGradient, feature occlusion, DeepLIFT, etc. Because time series are often z-normalized per sample to have a zero mean mainly for training stability, the zero value does not correspond to the lack of feature even if the original unnormalized time series does. For example, power consumption time series (e.g. ElectricDevices and LargeKitchenAppliances datasets in UCR repository) or time series dataset containing number of events or magnitude of events (e.g. Earthquakes and ChinaTown datasets in UCR repository or network traffic classification [19]) are examples of natural time series where zero corresponds to a lack of feature/event in the unnormalized version, but not in the normalized one. This causes many artifacts to appear in XAI methods as shown in Figure 2a. This is consistent with the observation in [20] that a global mean shift in data changes the explanation.
@@
where $x_{min}$ is the minimum value in the normalized time series. Note that the reverse operation ($F^{-1}$) is needed to be embedded into a new neural network, $M(F^{-1}(\cdot))$. This can be easily done by adding a single fully connected layer before $M()$ where it adds the last element to the first N elements. As shown in Figure 2b, generating the attribution on mean zero space using the same method as Figure 2a completely removes all artifacts assigned to the part with no signal.
@@
Difference Space is a result of taking a difference of each two consecutive time steps as a new time series. Differencing has been widely used in time series analysis to remove a trend or to make a time series stationary [21]. In our framework, difference domain has two benefits. First, for non-stationary time series (often with trend), such as stock price or atmospheric time series, a shapelet-like feature in a small region can be completely imperceptible. For example, in a synthetic data generated to mimic a non-stationary time series with a small shapelet as a distinguishable feature, a shapelet feature may not be easily visible when it is small in comparison to the range of the entire time series [^3] We use synthetic data in this experiment since there is no dataset with such properties in UCR repository., as shown in Figure 3a. However, the shapelets (i.e. semi-linear increase/decrease patterns visible in difference domain) can be clearly located in difference domain where the long range fluctuations is effectively removed, as shown in the corresponding samples in Figure 3b. Interestingly, the difference and time domain has a huge impact on the performance of XAI methods. The XAI method completely fails the time domain which can be explained by the second benefit.
Second, the difference domain implicitly changes the notion of a feature. Here, the lack of change in time domain corresponds to zeros in difference domain (implying a lack of feature), and changes correspond to non-zeros values. This can potentially solve the baseline issue for certain types of time series. For example, in time series associated with movements (e.g. GunPoint) or ECG signals, the lack of activity is manifested by a sequence of near constant values, not necessarily zeros. By taking the difference of the time series, a sequence of unchanging values is mapped to a sequence of zeros and, hence, aligns with the most explanation methods’ baselines. Formally, the difference space is defined as
@@
The reverse operation is trivial. Figure 2c and 2d shows the attribution generated by DeepLIFT on time domain and difference domain, respectively. Note that in Figure 2d we show the attribution generated on difference space on top of the time domain just for better visualization. In reality, the difference domain time series looks different from the original time domain. As shown, the difference space’s explanation is sparser and focuses on the actual heartbeat of the ECG signal.
(a) Original TS
(b) Decomposed component 1
(c) Decomposed component 2
Figure 4: DeepLIFT attribution on a synthetic dataset for the original TS versus the calibrated attribution on the decomposed components.
Decomposition Space is a process of generating explanation on each component of the decomposed time series. Time series decomposition is a widely used method in many fields, such as financial time series [22], weather forecasting [23], traffic prediction [24], etc. Although deep models are mostly trained non-decomposed time series, attribution methods on the non-decomposed time series can be hard to interpret. For example, seasonality and trends can easily obscure the shapelet pattern. To solve this issue, we use the singular-spectrum analysis (SSA) decomposition method to get a set of time series with different components. Then, the reverse operation adds these components to reconstruct the original time series. Hence, we can generate an attribution for each of these components separately. In Figure 4, we synthesize a dataset with different trends and multiple seasonality cycles plus some class-specific shapelet features. The ResNet model achieves $100\%$ accuracy when trained on the aggregated time series. As shown in Figure 4a, DeepLIFT correctly highlights the location of the shapelet. However, because of the trend and seasonality, it is hard to identify the shapelet’s pattern. The second decomposed component (Figure 4c) containing the zigzag shapelet can be easily interpreted.
There is one caveat when using back-propagation methods for decomposition time series: Because the reverse operation simply adds the decomposed components, the back-propagation methods essentially distribute the attribution similarly to all components. To solve this issue, we calibrate the attribution by combining the occlusion method with back-propagation methods as follows: Each decomposed component is occluded separately and the confidence drop is measured. Then, the confidence drop of each component is multiplied by the attribution values of that component. Finally, all attributions are normalized.
IV Evaluation
@@
Experimental Setup: We include several datasets covering different types of time series, as shown in Table I, mainly from UCR Repository [25] except AudioMNIST provided in [15]. For each dataset, we trained two commonly used model architectures: a ResNet [^4] ResNet implementation for time series is adopted from TSEvo package [26]. and an InceptionTime model [27]. We emphasize that all models are trained only on time domain. The explanations are generated on different domain using the process described in Method Section. We evaluate nine well-known XAI methods: DeepLIFT, GradientSHAP, Guided BackProp, InputXGradient (I×$\times$×G), Integrated Gradient (IG), KernelSHAP, LIME [28], Occlusion, and Saliency [5]. We use Captum implementation [29] for XAI methods. For min zero space, we ignore the attribution of the last time step, which is a placeholder for the min value. The reason is that perturbation of this single point, unlike other time points, changes the entire time series when the reverse operation is done. This causes the majority of XAI methods to give this point a very large attribution. Furthermore, none of the existing datasets in the UCR repository is suitable for the decomposition space since they are mostly short with no trends or seasonality.
@@
Note that the goal of this paper is not to have comprehensive evaluation of different XAI methods, as it has previously shown that no single XAI method can outperform others in all metrics/datasets [30]. Rather, we focus on illustrating that the concept of spaces is useful and that the same XAI method may bring a better explanation on different spaces depending on the type of time series.
@@
TABLE I: Dataset types and models’ accuracy
IV-A Evaluation Criteria
Robustness: Robustness of an attribution method refers to its sensitivity with respect to small perturbations of the target sample [18]. Unlike previous studies, we investigate the effect of different input spaces. As a result, a trained model wrapped to take frequency domain input may be more sensitive than the unwrapped version taking the time domain input. It is not trivial whether a model wrapped in other domains is more sensitive to perturbation than time domain or not. Hence, we also need to evaluate the sensitivity of the model itself in different input spaces. So, unlike previous work, we define two robustness: the first is associated with the classifier and the second is associated with the XAI method. Classifier robustness is defined as follow: